OpenStudy (anonymous):

the value of 2^(1/4).4^(1/8).8^(1/16).16(1/32)........................ is?

OpenStudy (anonymous):

2^(1/4+2/8+3/16+.....)

OpenStudy (anonymous):

find the sum of AGP series....1/4+2/8+3/16+4/32

OpenStudy (anonymous):

2^(13/16)

OpenStudy (anonymous):

this is not an AP, the 1st n 2nd term are the same isnt it?

OpenStudy (anonymous):

yes it is AGP....

OpenStudy (anonymous):

1/2^2 +(1+1)/2^3+(1+2)/2^4+(1+3)/2^5

OpenStudy (anonymous):

AGP is airthmatic geomatric series

OpenStudy (anonymous):

got it?????

OpenStudy (anonymous):

tn=n/2^(n+1)

OpenStudy (anonymous):

from tn you have to find Sn. here the series is small so you can add but otherwise you have to use series method.........

OpenStudy (anonymous):

whats the series method??

OpenStudy (anonymous):

check the page http://www.ilovemaths.com/3sequence.asp

OpenStudy (anonymous):

this is an infinite series i guess we gotta use Sn= a/1-r need to break the question itno that form

OpenStudy (anonymous):

no this form will not be applicable........

OpenStudy (anonymous):

tn=n/2^(n+1) \[t _{n+1}=(n+1)/2^{n+2}=n/2^{n+2}+1/2^{n+2}\] \[t _{n+1}=t _{n}/2+1/2^{n+2}\] t2=t1/2+1/2^3 t3=t2/2+1/2^4 . . . --------------------- adding Sn-t1=Sn/2 +(1/2^3+1/2^4+.........) Sn/2=t1+(1/2^3+1/2^4....)

OpenStudy (anonymous):

Sn/2=1/4+(1/2^3)(1/(1-1/2)) =1/4+1/4 Sn=1

OpenStudy (anonymous):

hence 2^1=2

OpenStudy (anonymous):

Got it?????

OpenStudy (anonymous):

i think u have understood